OpenStudy (anonymous):

Write an expression using rational Numbers to represent the problem, then calculate... Can anyone show me how I should do this?

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

mark has 24 news papers to deliver. in one appartment building he delivers 3/8 of the,. in the next appt he delivers 2/3 of the remaining amount . how many does he have left to deliver... i can figure the qustion out but i am not sure how to show it on my paper

OpenStudy (lgbasallote):

3/8 of the what?

OpenStudy (anonymous):

news papers

OpenStudy (lgbasallote):

in the first apartment he delivers 3/8 of the newspapers so \[24 \times \frac 38\] so what's left of that is \[24 \times \frac 58\] he delivers 2/3 of that to the second apartment so.. \[24 \times \frac 58 \times \frac 23\] so the remaining would be \[24 \times \frac 58 \times \frac 13\] got it?